A sample of 100 items is taken at random from a batch known to contain 40%
defectives. What is probability that the sample contains: (i) at least 44 defectives,
(ii) exactly 44 defectives?
1
Expert's answer
2020-03-09T13:27:25-0400

It is a Binomial problem with"P=0.4, n=100"

i. "P(\ge44)"

"P(\ge44)=1- P(\le43)"

From excel Binomial formula; =BINOM.DIST(43,100,0.4,TRUE),"P(\le43)=0.763469"

Thus, "P(\ge44)=1-0.763469=0.236531"

ii. "P(=44)"

"P(=44)" is given by =BINOM.DIST(44,100,0.4,FALSE) Excel formula, which yields 0.057629547.

Thus,

"P(=44)= 0.057629547"